Two cars, which are 612 miles apart and whose speeds differ by 10 mph, are moving toward each other. If they meet in 6 hours, find the speed of each car.
Answer by rfer(16322)

612/6=102 in one hr
x+x+10=102
2x=92
x=92/2
x=46
x+10=56